# Adaptive Hyper-Graph Convolution Network for Skeleton-based Human Action Recognition with Virtual Connections

### v1

@@ III-E Virtual Connections
It is worth emphasising that incorporating learnable joints among different samples is essential for enhancing the model capacity, as these learnable joints capture generalised features of human actions. This not only enriches the semantic information but also facilitates easier interaction connections among real joints. Therefore, we introduce the hyper-joints which are to participate in the hyper-graph convolution as shown in Figure 2 (a). The shape of hyper-joints is consistent with the feature $\mathbf{F_{p}}\in\mathbb{R}^{C^{{}^{\prime}}\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C^{{}^{\prime}}\times T\times V_{h}}$. Specifically, the learning of hyper-joints is supervised by the loss function, which guides these hyper-joints to support generalisable driven features embedded within large amounts of data. Furthermore, we set independent hyper-joints at each layer of Hyper-GCN, aiming to harmonise features at different depths. Typically, these hyper-joints are involved in spatial hyper-graph convolution rather than in temporal convolution.
@@
$C=\frac{\text{ReLU}(\mathbf{F_{h}F^{T}_{h}})}{||\mathbf{F_{h}}||^{2}}-\mathbf{I},$ (12)
where $\mathbf{I}$ represents the identity matrix which is used to emphasise the irrelevance between each learned hyper-joint pair.
Specifically, in Divergence Loss, we measure the differences of hyper-joints with the mean of the cosine matrix $C$ in each layer. The calculation can be formulated as:
$\mathcal{L}_{h}(C)$ $=\frac{\sum^{V_{h}}_{i=1}\sum^{V_{h}}_{j=1}c_{i,j}}{(V_{h}-1)^{2}},\quad c_{i,j}\in C,$ (13)
@@
where $V_{h}$ represents the number of hyper-joints we introduced. $C_{l}$ represents the cosine matrix of the $l$-th layer. Additionally, to preserve the original physical topology, we use the AHC-Module solely to learn the topology between feature $\mathbf{F_{h}}$ and feature $\mathbf{F_{p}}$.
@@ III-F Entire Architecture
Hyper-GCN consists of an embedding layer, 9 spatial-temporal convolution layers, a global average pooling, and an FC layer, as shown in Figure 2 (c).
The embedding layer is utilised to map individual joints into a high-dimensional feature space for subsequent adaptive hyper-graph convolution. To preserve the position awareness of the embedded features, we introduce the position embedding in the embedding layer inspired by InfoGCN [12].
The 9 layers of Hyper-GCN are categorised into 3 stages. Each stage consists of 3 layers. Each layer consists of Multi-Scale Hyper-graph convolution (MS-HGC) and Multi-Scale Temporal convolution (MS-TC) [13]. In stage 1, the number of channels is set as 128. Then it increases to 256 for stage 2 and stage 3.
Dense connections are introduced to integrate deep and shallow features in each stage, which can smooth the obtained distribution against the hyper-vertices interaction at each layer.
Methods Publication Modalities GFLOPs Params (M) NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
ST-GCN [18] AAAI 2018 J+B - - 81.5 88.3 70.7 73.2 -
2s-AGCN [20] CVPR 2019 J+B - - 88.5 95.1 82.5 84.2 -
DGNN [32] CVPR 2019 J+B+JM+BM - - 89.9 96.1 - - -
SGN [30] CVPR 2020 J+B+JM+BM - - 89.0 94.5 79.2 81.5 92.5
DC-GCN+ADG [33] ECCV 2020 J+B+JM+BM 1.83 4.9 90.8 96.6 86.5 88.1 95.3
DDGCN [31] ECCV 2020 J+B+JM+BM - - 91.1 97.1 - - -
MS-G3D [42] CVPR 2020 J+B+JM+BM 5.22 2.8 91.5 96.2 86.9 88.4 -
MST-GCN [34] AAAI 2021 J+B+JM+BM - 12.0 91.5 96.6 87.5 88.8 -
CTR-GCN [13] ICCV 2021 J+B+JM+BM 1.97 1.5 92.4 96.4 88.9 90.6 96.5
EfficientGCN-B4 [43] TPAMI 2022 J+B+JM+BM 15.2 2.0 91.7 95.7 88.3 89.1 -
InfoGCN* [12] CVPR 2022 J+B+J’+B’ 1.84 1.6 92.7 96.9 89.4 90.7 96.6
FR Head [48] CVPR 2023 J+B+JM+BM - 2.0 92.8 96.8 89.5 90.9 96.8
HD-GCN* [19] ICCV 2023 J+B+J’+B’ 1.77 1.7 93.0 97.0 89.8 91.2 96.9
DS-GCN [14] AAAI 2024 J+B+JM+BM - - 93.1 97.5 89.2 90.3 -
Hyper-GCN (w/o ensemble) - J 1.62 1.0 91.1 95.3 87.0 88.4 94.8
Hyper-GCN (2-ensemble) - J+B 1.62 1.0 92.9 96.3 89.8 90.9 96.2
Hyper-GCN (4-ensemble) - J+B+JM+BM 1.62 1.0 93.1 96.7 90.2 91.4 96.6
TABLE I: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets. For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods. J, B, JB, and JM represent the joint, bone, joint motion, and bone motion. In addition, for InfoGCN, HD-GCN [12, 19], we use the results from 4-streams according to the original papers for a fair comparison. GFLOPs and Params (M) all represent the single-stream.
@@ IV Evaluation
In this section, we conduct extensive experiments on standard benchmarking datasets to evaluate our Hyper-GCN, in the task of skeleton-based human action recognition.
@@ IV-A Datasets
NTU-RGB+D 60 NTU-RGB+D 60 [11] is a large dataset widely used in skeleton-based human action recognition. It consists of 56,880 different samples, categorised into 60 classes, obtained from the directed performances of 40 different actors. There are two 2 evaluation benchmarks. (1) Cross-Subject (X-Sub): For the 40 actors, 20 are used for training and 20 for validation. (2) Cross-view (X-View): For 3 views, 2 are used for training, and 1 for validation.
NTU-RGB+D 120 NTU-RGB+D 120 [25] is am extention of NTU 60, which introduce 57,367 new action samples. It consists of 114,480 different samples, categorised into 120 classes instead of 60 classes. The number of actors increase to 106. It also corresponds yo 2 benchmarks. (1) Cross-Subject (X-Sub): For the 106 actors, 53 are used for training and 53 for validation. (2) Cross-Setup (X-Set): For the 32 setups, the sample with even setup IDs are used for training, and the odd setup IDs for validation.
Northwestern-UCLA. The Northwestern-UCLA (NW-UCLA) dataset [49] contains 1494 video clips, which is categorised into 10 classes. It is obtained through 3 cameras with different camera views and is performed by 10 actors. As the proposed benchmark, for 3 camera views, 2 camera views are used for training, and the remaining 1 is used for validation.
@@ IV-B Implementation Details
Our training and evaluation stages are on a single RTX 3090 GPU. In terms of the training phase, Hyper-GCN is optimised by Stochastic Gradient Descent (SGD) with Nesterov momentum set at 0.9 and a weight decay at 0.0004. Our implementation uses label smooth cross-entropy loss with the Divergence Loss we proposed. We set a total of 140 epochs with the start 5 warm-up epochs. The initial learning rate is 0.05, which is reduced to 0.005 at epoch 110 and to 0.0005 at epoch 120. The training batch size is set to 64 for NTU RGB+D and NTU RGB+D 120, and 16 for Northwestern-UCLA.
@@ IV-C Comparison with the State-of-the-Art
Multi-stream ensembles proposed in  [50] have been proven effective by most of the existing state-of-the-art methods. We also use a 4-streams ensemble to evaluate our model.
The detailed results are reported in Table I. To the best of our knowledge, a comparison with SOTA shows that our proposed Hyper-GCN achieves the SOTA level on each dataset. Especially, we outperform the involved competitors on NTU-RGB+D 120. This demonstrates the effectiveness and merits of Hyper-GCN.
@@ IV-D Ablation Study
We conduct ablation experiments and report visualisations to analyse the effectiveness of each component of our design. All the ablation study is evaluated on the benchmark of NTU-RGB+D 120, using the joint modality. The Baseline for comparison is set by replacing the MS-HGC with vanilla GC, and removing the hyper-joints and dense connections in the backbone.
Figure 4: Top-1 accuracy in selected classes. The $K$-scale represents the $K$-scale hyper-graph in MS-HGC. The data is displayed starting from 75%.
$K$-scale
NTU RGB+D 120
X-Sub (%)
Baseline - 85.1
Single scale 3 86.4 ($\uparrow$ 1.3)
5 86.5 ($\uparrow$ 1.4)
7 86.5 ($\uparrow$ 1.4)
9 86.3 ($\uparrow$ 1.2)
Multi scale [0,2,3,4,5,6,7,8] 86.6 ($\uparrow$ 1.5)
[2,3,4,5,6,7,8,9] 86.7 ($\uparrow$ 1.6)
[3,4,5,6,7,8,9,10] 86.6 ($\uparrow$ 1.5)
TABLE II: Comparison of the hyper-parameter $K$ in terms of adaptive hyper-graph construction. [$k_{1}$,$k_{2}$,…,$k_{8}$] represents the $K$-scale hyper-graph of 8 branches in MS-HGC. The bold font represents the best results.
@@
Multi-Scale Hyper-graph: As we mentioned, Multi-Scale Hyper-graph Convolution (MS-HGC) consists of an 8-branch parallel architecture. In particular, each branch corresponds to a hyper-graph at a different scale. The hyper-graphs at different scales represent the maximum number of hyper-edges each joint is contained in. They reflect the varying degrees of aggregation in the constructed topology. The different scales of hyper-graphs are determined by the hyper-parameter $K$ that we define. Therefore, to explore how to set the hyper-parameter $K$ for the branches in MS-HGC, we design the ablation experiments to investigate the impact of different scales on the model’s recognition ability, as well as to compare the results by defining a single scale.
@@
The results are reported in Table II. When $K$ equals 1, each hyper-edge contains only one joint, which essentially eliminates any interaction between joints. Therefore, we do not set K to 1. Observing the design of the single scale hyper-graph, it becomes clear that increasing $K$ does not always yield better performance. A very large $K$ means that the maximum number of hyper-edges each joint is contained in increases, which in turn causes the number of joints in each hyper-edge to grow. Even though these ”extra” joints may be optimised for minimal weight, they are likely to act as ”noise”, disrupting the information interaction represented by the hyper-edge.
Furthermore, compared to a single-scale design, it is evident that the multi-scale design of MS-HGC outperforms the single scale design. For the complex and diverse characteristics of human action, multi-scale hyper-graphs in the spatial dimension can better capture features at multiple levels. Therefore, we visualise the Top-1 accuracy of selected human action categories as shown in Figure 4. This is especially advantageous for recognising actions with varying degrees of motion. For instance, for actions like ”Brushing teeth” or ”Salutation”, where only local joints are in motion, the smaller-scale hyper-graphs focus more on the local joints. In contrast, for actions like ”Stand up” and ”Jump up”, where the entire body is engaged, a larger-scale hyper-graph can capture information between distant joints. This also inspired us to adopt the AHC-Module in different branches to complementarily learn hyper-graph topology at different scales.
NTU RGB+D 120
X-Sub (%)
GFLOPs Parmas (M)
Baseline 85.1 1.70 1.52
w/ MS-HGC 86.6 ($\uparrow$ 1.5) 1.59 1.03
w/ 1 86.7 ($\uparrow$ 1.6) 1.61 1.03
w/ 2 86.9 ($\uparrow$ 1.8) 1.62 1.04
w/ 3 87.0 ($\uparrow$ 1.9) 1.62 1.05
w/ 4 86.6 ($\uparrow$ 1.5) 1.64 1.05
w/ 5 86.4 ($\uparrow$ 1.3) 1.66 1.06
TABLE III: The comparison of the configurations of hyper joints. w/ $N$ represents the number of hyper joints in each layer. These experiments are conducted in the setting w/ MS-HGC. The bold font represents the best results.
@@
Figure 5: The comparison of the cosine matrices of hyper-joints. The matrix elements range from $[-1,1]$. This experiment is set w/ 3 hyper-joints. Part (a) represents w/o Divergence Loss. Part (b) represents w/ Divergence Loss. The cosine matrix is calculated by Eqn. 12.
@@
Virtual Connections: We conduct the ablation study to analyse the hyper joints with the Divergence Loss. The results are listed in Table III. By comparison, the performance is not monotonically increasing with the number of hyper joints. The case of introducing only 3 hyper joints achieves the best performance. As our hyper joints are learned from a large amount of data, too many hyper joints can introduce redundant and ambiguous clues, weakening the model capacity. Besides, the involvement of hyper joints delivers consistent improvement, while introducing only a marginal increase in terms of model size and complexity.
@@
In addition, we visualised the cosine matrices in certain layers of Hyper-GCN, as shown in Figure 5. Clearly, in the absence of the Divergence Loss, the homogenisation of the learned hyper-joints in several layers is severe. This can limit the ability of hyper-joints to represent the generalised features of human actions. The observation further validates the effectiveness of the Divergence Loss for hyper-joint optimisation, supporting our viewpoint that generalised features with a certain degree of differentiation are needed to participate in hyper-graph convolution.
Method S-1 S-2 S-3 MS-HGC H NTU-RGB+D 120
X-Sub
Baseline 84.8
$\checkmark$ 84.8
$\checkmark$ 84.8
$\checkmark$ 84.9 ($\uparrow$ 0.1)
Hyper-GCN $\checkmark$ $\checkmark$ 84.9 ($\uparrow$ 0.1)
$\checkmark$ $\checkmark$ 85.0 ($\uparrow$ 0.2)
$\checkmark$ $\checkmark$ $\checkmark$ 85.1 ($\uparrow$ 0.3)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 86.6 ($\uparrow$ 2.0)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 87.0 ($\uparrow$ 2.4)
TABLE IV: The ablation study of architecture settings. S-$i$ represents adding dense connection in stage $i$. MS-HGC and H represent w/ MS-HGC and w/ 3 hyper-joints.
@@
Figure 6: t-SNE visualisation of Baseline and Hyper-GCN. Different colours represent the features of different joints.
Architecture of Backbone: We validate the effectiveness of performing dense connections in the backbone. The experimental results are collected in Table IV. The dense connections in the experiments are only added within each stage. It is observed that dense connections in shallow layers do not improve the performance as much as in deep layers. As the layers go deeper, the number of channels gradually increases and the variance of information extracted from each layer becomes steeper. This suggests that the deeper stages require dense connections to modulate the variance between features. For example, the performance introducing dense connections at stage 1 and stage 2 is close to adding the dense connection only at stage 3.
@@
To further validate that Hyper-GCN increases the efficiency of information interaction, we perform t-SNE projections on the output features of each layer, as shown in Figure 6. Notice that the semantic information represented by the last layer features of Hyper-GCN is very convergent. This represents that the semantic information of each joint is adequately conveyed in Hyper-GC. The joints with similar semantic can prevent sacrificing the information after global average pooling.

### v2

@@ III-D Multi-head Hyper-graph Convolution
Besides the hyper-graph, we incorporate the physical topology to emphasise the natural physical relations of a human being’s skeleton. To achieve this objective, existing methods [18, 27, 13, 19] divide the physical topology of the human body into 3 subsets. They are represented by $S=\{s_{id},s_{cf},s_{cp}\}$, where $s_{id}$, $s_{cf}$ and $s_{cp}$ denote the identity, centrifugal, and centripetal joint subsets. In our M-HGC, to ensure that the integrated physical topology in each head remains complete, we merge them into a single set. 8 heads are aggregated by channel concatenation after the hyper-graph convolutions, as shown in Figure 2 (a). The operation of M-HGC can be formulated as follows:
$\mathbf{F_{out}}$ $=\uplus^{8}_{i=1}{(\hat{A_{i}}+\alpha\cdot\hat{H_{i}})\mathbf{F^{k}_{in}}P_{i}},$ (11)
$\hat{A}$ and $\hat{H}$ denote the normalized physical adjacency matrix and normalized incidence matrix. $P$ is the learnable weight parameter for feature transform. $\alpha$ is a learnable parameter for the topology fusion of each head.
Category Methods Publication Modalities NTU-RGB+D 60 NTU-RGB+D 120 NW-UCLA (%)
X-Sub (%) X-View (%) X-Sub (%) X-Set (%)
GCN-based MST-GCN [40] AAAI21 J+B+JM+BM 91.5 96.6 87.5 88.8 -
CTR-GCN [13] ICCV21 J+B+JM+BM 92.4 96.4 88.9 90.6 96.5
EfficientGCN-B4 [49] TPAMI22 J+B+JM+BM 91.7 95.7 88.3 89.1 -
InfoGCN [12] CVPR22 J+B+JM+BM 92.7 96.9 89.4 90.7 96.6
FR Head [55] CVPR23 J+B+JM+BM 92.8 96.8 89.5 90.9 96.8
HD-GCN* [19] ICCV23 J+B+J’+B’ 93.0 97.0 89.8 91.2 96.9
DS-GCN [14] AAAI24 J+B+JM+BM 93.1 97.5 89.2 90.3 -
BlockGCN [20] CVPR24 J+B+JM+BM 93.1 97.0 90.3 91.5 96.9
HGCN-based Hyper-GNN [32] TIP21 J+B+JM+BM 89.5 95.7 - - -
Selective-HCN [33] ICMR22 J+B+JM+BM 90.8 96.6 - - -
DST-HCN [56] ICME23 J+B+JM+BM 92.3 96.8 88.8 90.7 96.6
Hyper-GCN (w/o ensemble) - J 91.4 95.5 87.0 88.7 95.5
Hyper-GCN (4-ensemble) - J+B+JM+BM 93.4 97.1 90.5 91.7 97.1
TABLE I: Comparison of Hyper-GCN with advanced solutions on NTU-RGB+D 60, NTU-RGB+D 120, and NW-UCLA datasets. For a fair comparison, we use a 4-streams ensemble to evaluate the performance of the mentioned methods. J, B, JB, and JM represent the joint, bone, joint motion, and bone motion. In addition, for HD-GCN [19], we use the results from 4-streams according to the original papers for a fair comparison. The bold font and the underline font represents the best result and 2-nd best result.
@@ III-E Virtual Connections
It is worth emphasising that incorporating learnable joints among different samples is essential for enhancing the model capacity, as these learnable joints capture generalised features of human actions. This not only enriches the semantic information but also facilitates easier interaction connections among real joints. Therefore, we introduce the hyper-joints which are to participate in the hyper-graph convolution as shown in Figure 2 (a).
The shape of hyper-joints is consistent with the feature $\mathbf{F_{p}}\in\mathbb{R}^{C\times T\times V}$, which is described as $\mathbf{F_{h}}\in\mathbb{R}^{C\times T\times V_{h}}$. Specifically, the learning of hyper-joints is supervised by the loss function, which guides these hyper-joints to support generalisable driven features embedded within large amounts of data. Furthermore, we set independent hyper-joints at each layer of Hyper-GCN, aiming to harmonise features at different depths. Typically, these hyper-joints are involved in spatial hyper-graph convolution rather than in temporal convolution.
@@
$C=\frac{\mathbf{F_{h}F^{T}_{h}}}{||\mathbf{F_{h}}||^{2}},$ (12)
Specifically, in Divergence Loss, we measure the differences of hyper-joints with the mean of the cosine matrix $C$ in each layer. Since the correlation of hyper-joints themselves cannot be optimised, we manually eliminate this part by minus the $V_{h}$. The calculation can be formulated as:
$\mathcal{L}_{h}(C)$ $=\frac{\sum^{V_{h}}_{i=1}\sum^{V_{h}}_{j=1}ReLU(c_{i,j})-V_{h}}{V_{h}(V_{h}-1)},\quad c_{i,j}\in C,$ (13)
@@
where $V_{h}$ represents the number of hyper-joints we introduced. $C_{l}$ represents the cosine matrix of the $l$-th layer.
Additionally, we manually connect the hyper joints to all the physical joints, which optimises the topology from the datasets.
@@ III-F Entire Architecture
Hyper-GCN consists of an embedding layer [12], 9 spatial-temporal convolution layers as shown in Figure 2 (c). Each layer consists of the proposed M-HGC and Multi-Scale Temporal convolution (MS-TC) [13]. 9 Layers are categorised into 3 stages with Dense connections. The channels in each stage are set to 128, 256 and 512.
@@ IV-A Datasets
NTU-RGB+D 60 & 120 NTU-RGB+D 60 [11] is a large dataset widely used in skeleton-based human action recognition, which is categorised into 60 classes. NTU-RGB+D 120 [34] is an extention to 120 classes of NTU 60. 4 benchmarks recommended by the official are adopted: (1) NTU60-XSub, (2) NTU60-XView, (3) NTU120-XSub, (4) NTU120-Xset.
Northwestern-UCLA. The Northwestern-UCLA (NW-UCLA) dataset [35] contains 1494 video clips, which is categorised into 10 classes. It contains 3 different camera views and is performed by 10 actors.
@@ IV-B Implementation Details
Our training and evaluation stages are on a single RTX 3090 GPU. In the training phase, Hyper-GCN is optimised by Stochastic Gradient Descent (SGD) with Nesterov momentum set at 0.9 and a weight decay at 0.0004. Our implementation uses label smooth cross-entropy loss with the Divergence Loss we proposed. We set a total of 140 epochs with the start 5 warm-up epochs. The initial learning rate is 0.05, which is reduced to 0.005 at epoch 110 and to 0.0005 at epoch 120.
@@ IV-C Comparison with the State-of-the-Art
Multi-stream ensembles proposed in  [57] have been proven effective by most of the existing state-of-the-art methods. We also use a 4-streams ensemble to evaluate our model. The detailed results are reported in Table I. To the best of our knowledge, a comparison with SOTA shows that our proposed Hyper-GCN achieves the SOTA level on each dataset. This demonstrates the effectiveness and merits of Hyper-GCN.
@@ IV-D Ablation Study
We conduct ablation experiments with visualisations to analyse the effectiveness of our design. All the ablation study is evaluated with the joint modality.
- NTU RGB+D 120 X-Sub (%)
K Uniform Non-uniform
Baseline - 84.7 84.7
w M-HGC 3 86.4 ($\uparrow$ 1.7) 86.0 ($\uparrow$ 1.3)
5 86.5 ($\uparrow$ 1.9) 86.2 ($\uparrow$ 1.5)
7 86.3 ($\uparrow$ 1.6) 86.5 ($\uparrow$ 1.8)
9 86.0 ($\uparrow$ 1.3) 86.7 ($\uparrow$ 2.2)
11 85.9 ($\uparrow$ 1.2) 86.4 ($\uparrow$ 1.7)
TABLE II: Ablation of the hyper-parameter $K$ in A-NHG. The Uniform represents K-uniform hyper-graph, in which one hyper-edge contains K joints. The non-uniform represents non-uniform hyper-graph, in which one joint is contained by $K$ hyper-edge.
Figure 4: Visualisation of hyper-graph in selected actions. The size of a joint means its weight within the hyper-edge. For ease of observation, we visualise the hyper-edges with the larger weight constructed in each sample.
@@
Hyper-graph Construction: The Non-uniform hyper-graphs with hyper-parameter $K$ represent the maximum number of hyper-edges contained in each joint. They reflect the varying degrees of aggregation in the constructed topology. In general, it is common to adopt a uniform hyper-graph with hyper-parameter $K$. Therefore, to explore how to set the hyper-parameter $K$ for the adaptive hyper-graph in M-HGC, we design the ablation experiments to investigate the impact of hyper-parameter $K$ on the model’s recognition ability, where both uniform and non-uniform hyper-graphs are considered.
@@
The results are reported in Table II. In both uniform and non-uniform hyper-graphs, a very large $K$ means that more joints are contained in each hyper-edge. Even though these ”extra” joints may be optimised for minimal weight, they are likely to act as ”noise”, disrupting the information interaction represented by the hyper-edge. As the $K$ decreases, it is difficult for the hyper-edge to represent complex joint combination patterns. Therefore, setting $K=5$ in uniform and $K=9$ in Non-uniform achieves the best performance. The non-uniform hyper-graph of A-NHG outperforms the uniform hyper-graph, as the combination of skeleton joints in different actions is diverse.
Furthermore, we visualise the hyper-graph inferred by Hyper-GCN on selected actions in Figure 4. The hyper-graph construction focuses on the motion joints that are most relevant to the corresponding action category. For “Kicking something” as part (d), the right leg joints and the left hand joint are modeled as a hyper-edge. Because in this action, the person’s right hand previously extends, and the right leg swings relatively to the right hand. Similarly, in the action ”Stand up” as shown in part (b), most joints of the upper body move relatively to right foot joint, which is not moving in the first half of frames. In addition, ”Taking a selfie” and ”Drinking water” are also consistent with the above view as Figure 4 (a), (c).
NTU RGB+D 120 X-Sub(%)
w/o $\mathcal{L}_{h}$ w/ $\mathcal{L}_{h}$
Baseline 84.7 84.7
w/o M-HGC w/ 1 84.9 ($\uparrow 0.2$) 84.9 ($\uparrow 0.2$)
w/ 3 84.9 ($\uparrow 0.2$) 85.2 ($\uparrow 0.5$)
w/ 5 84.7 85.0 ($\uparrow 0.3$)
w M-HGC w/ 1 86.7 ($\uparrow 2.0$) 86.7 ($\uparrow 2.0$)
w/ 3 86.6 ($\uparrow 1.8$) 86.9 ($\uparrow 2.2$)
w/ 5 86.6 ($\uparrow 1.8$) 84.8 ($\uparrow 2.1$)
TABLE III: The ablations on hyper joints. w/ $N$ represents the number of hyper joints in each layer. w/ $\mathcal{L}_{h}$ represents the training with Divergence Loss. The hyper-parameter $K$ of M-HGC is set to $9$.
@@
Figure 5: Comparison of the cosine matrices of hyper-joints. Part (a) and (b) represents w/o Divergence Loss and w/ Divergence Loss with 3 hyper-joints. It is calculated by Eqn. (12).
@@
Virtual Connections: We conduct the ablation study to analyse the hyper joints and the Divergence Loss. The results are listed in Table III. By comparison, the performance is not monotonically increasing with the number of hyper joints. The case of introducing only $3$ hyper joints achieves the best performance. As our hyper joints are learned from a large amount of data, involving a large number of hyper joints can introduce redundant and ambiguous clues, degrading the performance.
@@
In addition, we visualise the cosine matrices of Hyper-GCN, as shown in Figure 5. Clearly, in the absence of the Divergence Loss, the homogenisation of the learned hyper-joints is severe. This impedes the ability of hyper-joints to represent the generalised features of human actions. The observation further validates the effectiveness of the Divergence Loss for hyper-joint optimisation, supporting our viewpoint that generalised features with a certain degree of differentiation are needed to participate in hyper-graph convolution.
@@
Figure 6: t-SNE visualisation of Baseline and Hyper-GCN. Different colours represent different joints. Part (a) and (b) reprsent the output features in Baseline and Hyper-GCN.
@@
Effectiveness of Hyper-GCN: To further validate that Hyper-GCN increases the efficiency of information interaction, we perform t-SNE projections on the output features in last layer between Baseline and Hyper-GCN, as shown in Figure 6 for comparison. Notice that the semantic information represented by the last Hyper-GCN layer is very convergent. This represents that the semantic information of each joint is adequately conveyed in M-HGC. The joints with similar semantics can prevent sacrificing the information after global average pooling.
Methods Publications NTU RGB+D 120 GFLOPs Params
X-Sub(%) X-Set(%)
CTR-GCN[13] ICCV21 88.9 90.4 1.97 1.46
InfoGCN[12] CVPR23 89.4 90.4 1.70 1.52
FR-Head[55] CVPR23 89.5 90.9 2.0 -
BlockGCN[20] CVPR24 90.3 91.5 1.63 1.31
Hyper-GCN (-) - 90.4 91.5 1.63 1.06
DSTA-Net[25] ACCV20 86.6 89.0 16.18 3.45
IIP-Transformer[26] BigData23 89.9 90.9 7.20 2.90
SkateFormer[44] ECCV24 89.8 91.4 3.62 2.03
Hyper-GCN - 90.5 91.7 2.88 2.34
TABLE IV: The comparison of the performance and complexity. Hyper-GCN (-) represents the lightweight version.
IV-E Comparison of Complexity
Increasing the number of channels provides considerable improvement, which also increases the complexity to some extent. Given that advanced lightweight GCN-based methods are between 1M and 2M. For a fair comparison, we set channels in the last 3 layers to 256 to reduce the parameters to 1.06M, which is lighter weight than all involved counterparts. The performance of our lightweight version compared with GCN-based methods is reported in Table IV. In fact, on NTU RGB+D 120, which is the most challenge dataset, the performance of Hyper-GCN is degraded, but still achieves the SOTA level with the minimal parameter volume. In addition, on a comparable basis with the parameters of Transformer-based methods, our Hyper-GCN also outperform the SOTA.